# Explorer: Scaling Exploration-driven Web Trajectory Synthesis for Multimodal Web Agents

### v2

@@ Abstract
Recent success in large multimodal models (LMMs) has sparked promising applications of agents capable of autonomously completing complex web tasks. While open-source LMM agents have made significant advances in offline evaluation benchmarks, their performance still falls substantially short of human-level capabilities in more realistic online settings. A key bottleneck is the lack of diverse and large-scale trajectory-level datasets across various domains, which are expensive to collect. In this paper, we address this challenge by developing a scalable recipe to synthesize the largest and most diverse trajectory-level dataset to date, containing over 94K successful multimodal web trajectories, spanning 49K unique URLs, 720K screenshots, and 33M web elements. In particular, we leverage extensive web exploration and refinement to obtain diverse task intents. The average cost is 28 cents per successful trajectory, making it affordable to a wide range of users in the community. Leveraging this dataset, we train Explorer, a multimodal web agent, and demonstrate strong performance on both offline and online web agent benchmarks such as Mind2Web-Live, Multimodal-Mind2Web, and MiniWob++. Additionally, our experiments highlight data scaling as a key driver for improving web agent capabilities. We hope this study makes state-of-the-art LMM-based agent research at a larger scale more accessible.
@@ 1 Introduction
Graphical User Interfaces (GUIs) serve as the primary medium for user interaction across digital environments. Within the GUI environment, LLM-based agents Su et al. (2024) have shown great potential in automating complex workflows for human users. These agents are designed to operate across diverse interfaces, including the web Deng et al. (2023); Zhou et al. (2024); Zheng et al. (2024), desktop Xie et al. (2024); Wu et al. (2024); Bonatti et al. (2024), and mobile platforms Rawles et al. (2023); Yan et al. (2023). Navigating modern GUI interfaces, which integrate textual, graphical, and interactive components, typically requires agents to possess visual grounding, long-term planning, and memory management capabilities.
@@
Recent work Cheng et al. (2024); Gou et al. (2025) has demonstrated the effectiveness of synthetic data for enhancing visual grounding Gou et al. (2025); Chen et al. (2024a); Kapoor et al. (2024); Chen et al. (2024b) and planning Xu et al. (2024b); Zhang et al. (2024). Developing end-to-end GUI agents with long-term planning and grounding capabilities requires training on multi-step trajectory data Xu et al. (2025, 2024b); Qin et al. (2025). However, existing trajectory datasets are primarily human-annotated Deng et al. (2023); Li et al. (2024); Lu et al. (2024) or leverage synthetic data just for task proposal curation Lai et al. (2024); Chen et al. (2024a). And human annotation is expensive to scale for collecting large and diverse training datasets. Therefore, synthetic data has emerged as a promising alternative to human-annotated data Hartvigsen et al. (2022); Sahu et al. (2022); Ye et al. (2022); Tang et al. (2023); Mukherjee et al. (2023); Mitra et al. (2024). Collecting trajectory-level datasets presents unique challenges: 1) curating a diverse set of task intents at scale, 2) deploying an agent capable of interacting with a real-world environment to complete these tasks through a series of actions, and 3) verifying whether the task is accomplished by the executed action sequence.
@@
Figure 1: Data Generation Pipeline. The task proposer agent generates an initial task and the first action based on the website homepage. The task is then iteratively refined in subsequent steps by the refiner agent. Finally, the task summarizer agent constructs an overall task description from the action sequence, followed by task verification to assess correctness.
@@
Data diversity is essential for equipping generalist web agents with a broad range of skills. Existing work on synthetic web trajectory generation employs self-instruct for task proposal generation He et al. (2024b). It formulates task proposals from homepages or parametric LLM knowledge, overlooking the richer content available in deeper web pages, which is essential for achieving broader task diversity. Another line of work leverages web tutorials as a form of supervision for generating web trajectories Ou et al. (2024); Xu et al. (2025). While web tutorials effectively cover common daily user tasks, the resulting trajectory data exhibits limited domain diversity in terms of website and domain coverage (Table 1). Additionally, information-seeking tasks remain underrepresented. Due to these limitations, web agents trained on existing synthetic trajectory datasets have not seen much success in more realistic online evaluation settings. To enhance web agents’ performance in real-world settings, it is essential to incorporate greater diversity in their training trajectories.
@@
In this work, we develop a scalable and diverse web trajectory data synthesis recipe for training GUI agent models. Inspired by how humans learn to use the internet, we leverage exploration as a key mechanism for achieving diversity in task intents. We introduce Explorer (EXPLoration-driven web trajectORy gEneratoR), a framework for systematic web exploration to generate diverse, high-quality trajectory datasets. Unlike prior work that relies on static task proposals, Explorer dynamically explores web environments to curate diverse, real-world tasks. This exploration-based approach ensures broader task coverage and better generalization to real-world scenarios. We instantiate this framework using popular URLs from several sources, such as Tranco Pochat et al. (2019) and similarweb.com as seeds. Our dataset comprises 94K diverse web trajectories spanning 49K unique URLs, making it the largest web trajectory dataset to date. Each trajectory is richly annotated with artifacts such as screenshots, raw and set-of-mark Yang et al. (2023) annotated versions, HTML, and the accessibility tree, enabling comprehensive web agent training. To construct this dataset, we develop a multi-agent pipeline that starts with an abstract task proposal and iteratively refines it into a more specific task through web exploration (Figure 1). Unlike previous approaches, our pipeline generates tasks better grounded in real-world websites, improving task relevance and diversity. To demonstrate the effectiveness of our dataset, we train small language models using just the synthetic data and outperform existing web agent baselines by a significant margin. The main contributions of this work are as follows:
@@ 2 Related Work
2.1 LLM-based Web Agents
Recent advances in multimodal language models have facilitated the development of web agents — autonomous systems designed to interact with real-world websites to perform everyday tasks Deng et al. (2023); Hong et al. (2024); Cheng et al. (2024); Zheng et al. (2024). Key challenges for web agents include long-term planning, visual grounding, and memory management. To improve long-context understanding, WebAgent Gur et al. (2024) utilizes multiple LLMs - one for planning, summarization, and grounded program synthesis. SeeAct Zheng et al. (2024) adopts a two-step procedure of planning followed by grounding at each step using GPT-4 to accomplish web agent tasks. Another line of work employs a vision-only approach to train a GUI grounding model that directly predicts pixel coordinates for executing GUI agent tasks Cheng et al. (2024); Kapoor et al. (2024); Gou et al. (2025). However, a significant bottleneck remains — the lack of large-scale, high-quality web trajectory data for training robust agents. Our work presents a new framework for synthesizing large-scale web trajectory data to train end-to-end web agents.
2.2 Web Agent Benchmarks and Datasets
Early benchmarks for web tasks such as MiniWob++ Liu et al. (2018) focused on testing low-level actions on simulated websites. However, these simulated websites fail to capture the complexity of the real-world web. Mind2Web Deng et al. (2023) introduces a trajectory-level dataset with 2K tasks across 137 real-world websites and 31 domains. However, it employs a static evaluation method that penalizes alternative valid execution paths. To overcome this limitation, follow-up work has explored alternative evaluation approaches, including functional correctness-based evaluation in WebArena Zhou et al. (2024) and key-node-based evaluation in Mind2Web-Live Pan et al. (2024b). Towards the goal of making web agents more capable of performing realistic tasks, GAIA Mialon et al. (2024) and AssistantBench Yoran et al. (2024) introduce benchmarks that include time-consuming information-seeking tasks. In this work, we develop Explorer, a multimodal web agent trained on our synthetic dataset, and showcase its strong performance across online and offline benchmarks, including Mind2Web-Live, Multimodal-Mind2Web, and MiniWob++.
@@ 2.2 Web Agent Benchmarks and Datasets
Analyze Tesla’s stock performance over various time periods on Yahoo Finance.
@@
2.3 Data Synthesis for Web Agents
Early efforts to acquire trajectory data for training web agents primarily relied on crowd-sourcing Deng et al. (2023); Lu et al. (2024). However, human annotation is cost-prohibitive, prompting the adoption of synthetic data generation approaches to facilitate large-scale data collection. AutoWebGLM Lai et al. (2024) and GUIAct Chen et al. (2024a) utilize LLMs to generate task proposals, which human experts subsequently annotate. OpenWebVoyager He et al. (2024b) employs a web agent to execute auto-generated task descriptions. However, since these task descriptions are generated using LLMs without exploring a website, they fail to capture the full diversity of possible tasks on that website. Another line of work, including Synatra Ou et al. (2024) and AgentTrek Xu et al. (2025), leverages web tutorials to guide web trajectory generation. Meanwhile, concurrent effort Murty et al. (2024) employs an exploration-based trajectory generation in WebArena’s sandbox, while our work focuses on more realistic web agent evaluation on live websites. To address diversity limitations in prior trajectory synthesis work, we design a bottom-up web trajectory synthesis pipeline that explores websites dynamically while maintaining a coherent high-level task intent.
@@ 3 Data Recipe
We design an automatic web trajectory synthesis pipeline that explores websites to generate diverse web trajectories. It utilizes Playwright***https://playwright.dev/ to execute actions and collect metadata from real-world websites, starting from an initial URL.For a 4K subset of trajectories, we instruct GPT-4o to navigate to the target website by formulating a Google search query based on the task description. The metadata includes screenshots, HTML, A11y tree, and actions in grounded and natural language forms. The action space is given in Table 3.
3.1 Website Selection
Action Type Description Count
click [$\mathtt{elem}$] Click on $\mathtt{elem}$. 415K
type [$\mathtt{elem}$] [$\mathtt{text}$] Type $\mathtt{text}$ 62K
goto [$\mathtt{url}$] Go to $\mathtt{url}$. 26K
search_google [$\mathtt{query}$] Search for $\mathtt{query}$ on Google. 4K
scroll [$\mathtt{up/down}$] Scroll up or down. 213K
Table 3: Action space for web navigation in Explorer.
@@ 3.1 Website Selection
Figure 2: Histogram of token distribution across trajectories in Explorer. The red vertical line shows the ${90}^{th}$ percentile of the distribution.
Model Cost per trajectory
Mind2Web Deng et al. (2023) \$$0.85$
AgentTrek Xu et al. (2025) \$$0.55$
Explorer \$$0.28$
Table 5: Cost comparison with other trajectory generation approaches.
@@
3.2 Data Generation Pipeline
We aim to develop a generalized pipeline for web exploration to collect diverse web trajectory data. To enhance diversity, we adopt a bottom-up approach, starting with low-level actions and progressively shaping them into high-level task descriptions while maintaining a coherent task intent. In the first step, the proposer agent generates an abstract task, which is refined to a more specific task through a refinement process. For instance, starting from the Amazon homepage, the initial task proposal might be “Find today’s deals on Amazon”, which is progressively refined into “Proceed to checkout for the Amazon eero Pro 6E mesh Wi-Fi router with 36% off” (Figure 1). Since the agents execute actions alongside the refinement process, the generated tasks respect real-world constraints, such as product availability, available color options, and other specifications, ensuring practical applicability. Our pipeline consists of the following LLM-powered agentsWe use GPT-4o as the agent backbone throughout the data generation process.:
Task Proposer.
Given a website homepage, including its screenshot and accessibility tree, the task proposer agent generates diverse initial tasks that could be performed on that website. The task descriptions at this stage are instructed to be high-level and abstract versions of the real-world tasks, which will be refined into more specific tasks in later stages. Along with generating the task proposal, the agent proposes and executes the first action toward completing that task. Furthermore, the agent is instructed to halt upon encountering robot detection mechanisms, CAPTCHA verification, login prompts, or payment requests.
Task Refiner.
The task refiner agent receives the initial task proposal or the refined task description from the previous step, along with the corresponding action history as inputs. It then predicts the next action consistent with the input task description and the updated refined task description while incorporating the complete action history. By iteratively refining the task description after each action, the agent ensures that the updated task remains aligned with the action history.
@@ 3.2 Data Generation Pipeline
Inspired by Pan et al. (2024a), the task verifier agent receives the task description and action history, serving as a critic to evaluate whether the trajectory successfully completes the specified task. In addition to the screenshots of the trajectory, it also receives a markdown representation of the last page. This ensures the verifier has the full context of the website’s final state, even when the viewport cannot capture all the content. Such automatic evaluation of web trajectories has been widely adopted in prior work Xu et al. (2025); He et al. (2024a); Koh et al. (2024). Figure 1 illustrates the above pipeline. The prompts for the above agents are given in Appendix D.
@@ 3.3 Dataset Analysis
Explorer comprises web trajectories spanning diverse domains, including services, entertainment, shopping, travel, and information, ensuring broad task diversity. Figure 3 visualizes the domain and subdomain distribution, highlighting the dataset’s rich diversity. Sample tasks from Explorer are presented in Table 2. To the best of our knowledge, Explorer with 94K trajectories is the largest web trajectory dataset of this scale. Table 1 shows a comparison with existing web agent datasets from the literature. The detailed statistics are given in Table 4. Figure 2 shows the histogram of token distribution across the dataset.
Beyond diversity, Explorer is also highly scalable and cost-efficient. Our approach achieves a cost of \$0.28 per successful trajectory, making it approx. 2×2$\times$2 × more cost-effective than AgentTrek Xu et al. (2025) (which incurs \$0.55 per trajectory) and significantly cheaper than human annotation (Table 5). Unlike human annotation, which requires training crowd workers and continuous quality monitoring, Explorer’s automated generation pipeline eliminates these bottlenecks, ensuring scalability with minimal overhead. By integrating diversity, scalability, and cost efficiency, Explorer sets a new benchmark for generating large-scale web trajectory datasets, making it a valuable resource for training generalist web agents.
@@ 4 Experiments
Multimodal-Mind2Web.
Mind2Web-Live.
Mind2Web-Live is a benchmark modified from the original Mind2Web dataset to test web agents on live websites rather than static trajectories. The benchmark evaluates performance using a key-node-based evaluation approach rather than using a golden action sequence, requiring valid trajectories to reach annotated “key nodes” across 104 test tasks in Mind2Web. Since Mind2Web-Live relies on real-world dynamic websites, it encounters robot detection such as reCAPTCHA, which hinders testing Xu et al. (2024b). To address this, we select a subset of 83 test set tasks that remain consistently accessible throughout our tests. Following Pan et al. (2024b), we report the average step success rate, completion rate, and full task success rate (SR) on the test set. The average step success rate and completion rate represent the proportion of completed key nodes, using macro and micro averages, respectively.
MiniWob++.
@@
Mistral-7B-Instr.-0.3 Jiang et al. (2023) $32.8$ $29.5$ $24.1$ $9.6$
Qwen2-72B-Instruct Bai et al. (2023) $40.9$ $15.4$
@@
Table 6: Results on Mind2Web-Live benchmark. Missing values are denoted by –. The results for GPT-4 and Mistral-7B have been reproduced on our Linux servers. The results for GPT-3.5 and Qwen2-72B-Instruct have been taken from Pan et al. (2024b). The full task success rate represents the successful completion of all key nodes for a given task. The average step success rate represents the proportion of completed key nodes, macro-averaged across tasks. The completion rate represents the proportion of completed key nodes, micro-averaged across tasks. Task SR (1) represents task SR with a tolerance of up to one error/key node.

### v3

@@ Abstract
Recent success in large multimodal models (LMMs) has sparked promising applications of agents capable of autonomously completing complex web tasks. While open-source LMM agents have made significant advances in offline evaluation benchmarks, their performance still falls substantially short of human-level capabilities in more realistic online settings. A key bottleneck is the lack of diverse and large-scale trajectory-level datasets across various domains, which are expensive to collect. In this paper, we address this challenge by developing a scalable recipe to synthesize the largest and most diverse trajectory-level dataset to date, containing over $94$K successful multimodal web trajectories, spanning $49$K unique URLs, $720$K screenshots, and $33$M web elements. In particular, we leverage extensive web exploration and refinement to obtain diverse task intents. The average cost is $28$ cents per successful trajectory, making it affordable to a wide range of users in the community. Leveraging this dataset, we train Explorer, a multimodal web agent, and demonstrate strong performance on both offline and online web agent benchmarks such as Mind2Web-Live, Multimodal-Mind2Web, and MiniWob++. Additionally, our experiments highlight data scaling as a key driver for improving web agent capabilities. We hope this study makes state-of-the-art LMM-based agent research at a larger scale more accessible. [^1] Project website: https://osu-nlp-group.github.io/Explorer/
@@ 1 Introduction
Graphical User Interfaces (GUIs) serve as the primary medium for user interaction across digital environments. Within the GUI environment, LLM-based agents Su et al. (2024) have shown great potential in automating complex workflows for human users. These agents are designed to operate across diverse interfaces, including the web Deng et al. (2023); Zhou et al. (2024); Zheng et al. (2024, 2025), desktop Xie et al. (2024); Wu et al. (2024), and mobile platforms Rawles et al. (2023); Yan et al. (2023). Navigating modern GUI interfaces, which integrate textual, graphical, and interactive components, typically requires agents to possess visual grounding, long-term planning, and memory management capabilities.
@@
Recent work Cheng et al. (2024); Gou et al. (2025) has demonstrated the effectiveness of synthetic data for enhancing visual grounding Gou et al. (2025); Chen et al. (2024a); Kapoor et al. (2024); Chen et al. (2024b) and planning Xu et al. (2025b); Zhang et al. (2024). Developing end-to-end GUI agents with long-term planning and grounding capabilities requires training on multi-step trajectory data Xu et al. (2025a, b); Qin et al. (2025). However, existing trajectory datasets are primarily human-annotated Deng et al. (2023); Li et al. (2024); Lu et al. (2024) or leverage synthetic data just for task proposal curation Lai et al. (2024); Chen et al. (2024a). And human annotation is expensive to scale for collecting large and diverse training datasets. Therefore, synthetic data has emerged as a promising alternative to human-annotated data Hartvigsen et al. (2022); Sahu et al. (2022); Ye et al. (2022); Tang et al. (2023); Mukherjee et al. (2023); Mitra et al. (2024). Collecting trajectory-level datasets presents unique challenges: 1) curating a diverse set of task intents at scale, 2) deploying an agent capable of interacting with a real-world environment to complete these tasks through a series of actions, and 3) verifying whether the task is accomplished by the executed action sequence.
@@
Figure 1: Data Generation Pipeline. The task proposer agent generates an abstract task proposal and the first action based on the website homepage. The task is then iteratively refined in subsequent steps by the refiner agent. Finally, the task summarizer agent constructs an overall task description from the action sequence, followed by task verification to assess correctness.
@@
Data diversity is essential for equipping generalist web agents with a broad range of skills. Existing work on synthetic web trajectory generation employs self-instruct for task proposal generation He et al. (2024b). It formulates task proposals from homepages or parametric LLM knowledge, overlooking the richer content available in deeper web pages, which is essential for achieving broader task diversity. Another line of work leverages web tutorials as a form of supervision for generating web trajectories Ou et al. (2024); Xu et al. (2025a). While web tutorials effectively cover common daily user tasks, the resulting trajectory data exhibits limited domain diversity in terms of website and domain coverage (Table 1). Additionally, information-seeking tasks remain underrepresented. Due to these limitations, web agents trained on existing synthetic trajectory datasets have not seen much success in more realistic online evaluation settings. To enhance web agents’ performance in real-world settings, it is essential to incorporate greater diversity in their training trajectories.
@@
In this work, we develop a scalable and diverse web trajectory data synthesis recipe for training GUI agent models. Inspired by how humans learn to use the internet, we leverage exploration as a key mechanism for achieving diversity in task intents. We introduce Explorer, a framework for systematic web exploration to generate diverse, high-quality trajectory datasets. Unlike prior work that relies on static task proposals, Explorer dynamically explores web environments to curate diverse, real-world tasks. This exploration-based approach ensures broader task coverage and better generalization to real-world scenarios. We instantiate this framework using popular URLs from several sources, such as Tranco Pochat et al. (2019) and similarweb.com as seeds. Our dataset comprises $94$K diverse web trajectories spanning $49$K unique URLs, making it the largest web trajectory dataset to date. Each trajectory is richly annotated with artifacts such as screenshots, raw and set-of-mark Yang et al. (2023) annotated versions, HTML, and the accessibility tree, enabling comprehensive web agent training. To construct this dataset, we develop a multi-agent pipeline that starts with an abstract task proposal and iteratively refines it into a more specific task through web exploration (Figure 1). Unlike previous approaches, our pipeline generates tasks better grounded in real-world websites, improving task relevance and diversity. To demonstrate the effectiveness of our dataset, we train small language models using just the synthetic data and outperform existing web agent baselines by a significant margin. The main contributions of this work are as follows:
@@ 2 Related Work
Recent advances in multimodal language models have facilitated the development of web agents — autonomous systems designed to interact with real-world websites to perform everyday tasks Deng et al. (2023); Hong et al. (2024); Cheng et al. (2024); Zheng et al. (2024, 2025); Xue et al. (2025). Early efforts to acquire trajectory data for training web agents primarily relied on crowd-sourcing Deng et al. (2023); Lu et al. (2024). However, due to the high cost of human annotation, recent work has adopted synthetic data generation for large-scale collection. AutoWebGLM Lai et al. (2024) and GUIAct Chen et al. (2024a) utilize LLMs to generate task proposals, which human experts subsequently annotate. OpenWebVoyager He et al. (2024b) employs a web agent to execute auto-generated task descriptions. However, since these task descriptions are generated using LLMs without exploring a website, they fail to capture the full diversity of possible tasks on that website. Another line of work, including Synatra Ou et al. (2024) and AgentTrek Xu et al. (2025a), leverages web tutorials to guide web trajectory generation. Meanwhile, concurrent effort Murty et al. (2024) employs an exploration-based trajectory generation in WebArena’s sandbox, while our work focuses on more realistic web agent evaluation on live websites. To address diversity limitations in prior trajectory synthesis work, we design a bottom-up web trajectory synthesis pipeline that explores websites dynamically while maintaining a coherent high-level task intent. We refer readers to Appendix F for further discussion.
@@ 3 Data Recipe
We design an automatic web trajectory synthesis pipeline that explores websites to generate diverse web trajectories. It utilizes Playwright [^2] https://playwright.dev/ to execute actions and collect metadata from real-world websites, starting from an initial URL. [^3] For a $4$K subset of trajectories, we instruct GPT-4o to navigate to the target website by formulating a Google search query based on the task description. The metadata includes screenshots, HTML, A11y tree, and actions in grounded and natural language forms.
3.1 Website Selection
3.2 Data Generation Pipeline
We aim to develop a generalized pipeline for web exploration to collect diverse web trajectory data. To enhance diversity, we adopt a bottom-up approach, starting with low-level actions and progressively shaping them into high-level task descriptions while maintaining a coherent task intent. In the first step, the proposer agent generates an abstract task, which is refined to a more specific task through a refinement process (Figure 1). Since the agents execute actions alongside the refinement process, the generated tasks respect real-world constraints, such as product availability, available color options, and other specifications, ensuring practical applicability. Our pipeline consists of the following LLM-powered agents [^4] We use GPT-4o as the agent backbone throughout the data generation process.:
Task Proposer.
Given a website homepage, including its screenshot and accessibility tree, the task proposer agent generates diverse initial tasks that could be performed on that website. The task descriptions at this stage are instructed to be high-level and abstract versions of the real-world tasks, which will be refined into more specific tasks in later stages. Along with generating the task proposal, the agent proposes and executes the first action toward completing that task. Furthermore, the agent is instructed to halt upon encountering robot detection such as CAPTCHA verification, login prompts, or payment requests.
Task Refiner.
The task refiner agent receives the initial task proposal or the refined task description from the previous step, along with the corresponding action history as input. It then predicts the next action consistent with the input task description and the updated, refined task description while incorporating the complete action history. By iteratively refining the task description after each action, the agent ensures that the updated task remains aligned with the action history.
@@ 3.2 Data Generation Pipeline
Inspired by Pan et al. (2024a), the task verifier agent receives the task description and action history, serving as a critic to evaluate whether the trajectory successfully completes the specified task. In addition to the screenshots of the trajectory, it also receives a markdown representation of the last page. This ensures the verifier has the full context of the website’s final state, even when the viewport cannot capture all the content. To ensure data quality, trajectories that are incoherent or misaligned with the high-level intent are discarded during this stage. Such automatic evaluation of web trajectories has been widely adopted in prior work Xu et al. (2025a); He et al. (2024a); Koh et al. (2024). Figure 1 illustrates the above pipeline. The prompts for the above agents are given in Appendix D.
@@ 3.3 Dataset Analysis
Explorer comprises web trajectories spanning diverse domains, including services, entertainment, shopping, travel, and information, ensuring broad task diversity. Sample tasks from Explorer are presented in Table 2. Figure 2 visualizes the domain and subdomain distribution, highlighting the dataset’s rich diversity. To the best of our knowledge, Explorer with $94$K trajectories is the largest web trajectory dataset of this scale. Table 1 shows a comparison with existing web agent datasets from the literature. The detailed statistics are given in Table 3. Beyond diversity, Explorer is also highly scalable and cost-efficient. Our approach achieves a cost of \$$0.28$ per successful trajectory, making it approximately 2×2$\times$2 × more cost-effective than AgentTrek Xu et al. (2025a) (which incurs \$$0.55$ per trajectory) and significantly cheaper than human annotation (Table 4). Unlike human annotation, which requires training crowd workers and continuous quality monitoring, Explorer’s automated generation pipeline eliminates these bottlenecks, ensuring scalability with minimal overhead. By combining diversity, scalability, and cost efficiency, Explorer sets a new benchmark for generating large-scale web trajectory datasets, making it an invaluable resource for training generalist GUI agents.
Model Cost per trajectory
Mind2Web Deng et al. (2023) \$$0.85$
AgentTrek Xu et al. (2025a) \$$0.55$
Explorer \$$0.28$
Table 4: Cost comparison with other approaches.
@@ 4 Experiments
Mistral-7B-Instr. Jiang et al. (2023) $32.8$ $29.5$ $24.1$ $9.6$
Qwen2-72B-Instr. Bai et al. (2023) $40.9$ $15.4$
@@
Table 5: Results on Mind2Web-Live benchmark. Missing values are denoted by –. The results for GPT-4 and Mistral-7B-Instruct have been reproduced on our Linux servers. The results for GPT-3.5 and Qwen2-72B-Instruct have been taken from Pan et al. (2024b). The full task success rate represents the successful completion of all key nodes for a given task. The average step success rate represents the proportion of completed key nodes, macro-averaged across tasks. The completion rate represents the proportion of completed key nodes, micro-averaged across tasks. Task SR (1) represents task SR with a tolerance of up to one error/key node.
Multimodal-Mind2Web.
Mind2Web-Live.
Mind2Web-Live is a benchmark modified from Mind2Web to test web agents on live websites rather than static trajectories. The benchmark evaluates performance using a key-node-based evaluation approach rather than using a golden action sequence, requiring valid trajectories to reach annotated “key nodes” across $104$ test tasks in Mind2Web. Since Mind2Web-Live relies on real-world dynamic websites, it encounters robot detection such as reCAPTCHA, which hinders testing Xu et al. (2025b). To address this, we select a subset of $83$ test set tasks that remain consistently accessible throughout our tests. Following Pan et al. (2024b), we report the average step success rate, completion rate, and full task success rate on the test set.
MiniWob++.
Model Train Data Cross-Task Cross-Website Cross-Domain Avg.
Ele. Acc Op. F1 Step SR Ele. Acc Op. F1 Step SR Ele. Acc Op. F1 Step SR
In-Context Learning
GPT-3.5 $19.4$ $59.2$ $16.8$ $14.9$ $56.5$ $14.1$ $25.2$ $57.9$ $24.1$ $18.3$
GPT-4 $40.8$ $63.1$ $32.3$ $30.2$ $61.0$ $27.0$ $35.4$ $61.9$ $29.7$ $29.7$
SeeAct Zheng et al. (2024) $46.4$ $73.4$ $40.2$ $38$ $67.8$ $32.4$ $42.4$ $69.3$ $36.8$ $36.5$
Supervised Fine-Tuning
EDGE-9.6B Chen et al. (2024b) Syn. + M2W $30.0$ $21.1$ $22.4$ $24.5$
MiniCPM-3.1B Chen et al. (2024a) Syn. + M2W $23.8$ $86.8$ $20.8$ $20.3$ $81.7$ $17.3$ $17.9$ $74.5$ $14.6$ $17.6$
ScribeAgent-32B Shen et al. (2024) Syn. traj. $38.0$ $52.9$ $35.6$ $34.1$ $52.7$ $32.5$ $39.4$ $54.7$ $37.3$ $35.1$
AgentTrek-7B Xu et al. (2025a) Syn. + M2W $60.8$ $88.9$ $55.7$ $57.6$ $88.1$ $51.4$ $56.0$ $87.5$ $52.6$ $53.2$
Explorer-4B M2W $48.1$ $88.0$ $44.8$ $49.1$ $87.2$ $45.0$ $46.9$ $87.7$ $44.6$ $44.8$
Explorer-4B Syn. + M2W $53.4$ $88.1$ $50.7$ $55.6$ $89.5$ $51.4$ $49.8$ $88.8$ $47.2$ $49.8$
Explorer-7B M2W $51.8$ $88.0$ $48.3$ $56.3$ $89.7$ $52.0$ $50.9$ $88.9$ $48.1$ $49.5$
Explorer-7B Syn. + M2W $56.5$ $90.3$ $53.2$ $60.5$ $90.7$ $56.7$ $55.7$ $90.4$ $53.0$ $54.3$